How long does it take an electron to complete a circular orbit at right angles to a magnetic field of .0001 T?
How long does it right angles to a ma take an electron to complete a circular orbit at gnetic field of o.0001 T? aht An electron moving
Answer #1

qvB = mv^2/r

=> v/r = Bq/m

Remember angular speed is:

ω = v/r = 2π/T

Where T is the period of a single orbit. Thus:

2π/T = B(q/m)

=> T = 2π/(B(q/m))

q/m = 1.76x10^11 C/kg for an electron, and B = 1x10^-4 T So we have the period:

T = 2π/(1x10^-4 x 1.76x10^11) = 3.57x10^-7 s
period.